OpenStudy (anonymous):

Factor this trinomial completely. -12x3 + 45x2 + 12x

OpenStudy (anonymous):

-12x3 + 45x2 + 12x =-3x(4x^2 -15x-4) =-3x(4x^2 -16x+x-4) =-3x(4x(x-4) +1(x-4)) =-3x(x-4)(4x+1)

OpenStudy (aravindg):

\[\huge \color{green}{\text{WELCOME TO OPENSTUDY!}}\]

OpenStudy (anonymous):

Thank you!